Calcium has a work function of 2.7 eV
(a) What is the function of calcium expressed in joules? (b) What is the threshold frequency for calcium? (c) What is the maximum wavelength that will cause emission from calcium? (e=1.6×10^-19 C, ℎ=6.6×10^-34 Js , c=3.0×10^8 m/s) .
1
Expert's answer
2020-11-10T07:08:01-0500

a)

The work function is

"W=2.7 eV=4.32 \times10^{-19} J"

b)

The threshold frequency is

"f=\frac{W}{h}=\frac{4.32 \times10^{-19}}{6.626 \times10^{-34}}=6.52 \times 10^{14} Hz"

c)

The maximum wavelength is

"\lambda =\frac{hc}{W} =\frac{(6.626 \times10^{-34})(3.0 \times10^{8})}{4.32 \times10^{-19}}=460 nm"